A charge of -8.76 μC is traveling at a speed of 8.85 × 106 m/s in a region of space where there is a magnetic field. The angle between the velocity of the charge and the field is 56.0o. A force of magnitude6.14 × 10-3 N acts on the charge. What is the magnitude of the magnetic field?
Answer #1

Force is given by

F = qv B sin x

6.14* 10^-3 = 8.76* 10^-6* 8.85* 10^6* B sin 56

B = 9.553* 10^-5 T

======

Comment before rate in case any doubt, will reply for sure.. goodluck